# Discriminating image representations with principal distortions

### v1

@@ Abstract
Image representations (artificial or biological) are often compared in terms of their global geometry; however, representations with similar global structure can have strikingly different local geometries. Here, we propose a framework for comparing a set of image representations in terms of their local geometries. We quantify the local geometry of a representation using the Fisher information matrix, a standard statistical tool for characterizing the sensitivity to local stimulus distortions, and use this as a substrate for a metric on the local geometry in the vicinity of a base image. This metric may then be used to optimally differentiate a set of models, by finding a pair of “principal distortions” that maximize the variance of the models under this metric. We use this framework to compare a set of simple models of the early visual system, identifying a novel set of image distortions that allow immediate comparison of the models by visual inspection. In a second example, we apply our method to a set of deep neural network models and reveal differences in the local geometry that arise due to architecture and training types. These examples highlight how our framework can be used to probe for informative differences in local sensitivities between complex computational models, and suggest how it could be used to compare model representations with human perception.
@@ 1 Introduction
Biological and artificial neural networks transform sensory stimuli into high-dimensional internal representations that support downstream tasks, and these representations are often described in terms of their neural population geometry (Chung & Abbott, 2021). This idea has led to a multitude of proposed measures of representational similarity (Kriegeskorte et al., 2008; Yamins & DiCarlo, 2016; Kornblith et al., 2019; Williams et al., 2021; Klabunde et al., 2023), and these measures are often used to compare representations within a computational model to the representations within a brain. However, despite differing in architectures and training procedures, many computational models of perceptual or neural responses are equally performant on these representational similarity measures (Schrimpf et al., 2018; Tuckute et al., 2023; Conwell et al., 2022). Are these models functionally interchangeable, or are the datasets and methods that are used to test them simply insufficient to reveal their differences?
Often the similarity between two representations is quantified by measuring alignment of the representations over a set of natural stimuli that are relatively far apart in stimulus space. In this way, these measures capture notions of global geometric similarity between representations. However, systems with similar global structure can have strikingly different local geometries. For example, Szegedy et al. (2013) found image distortions that were imperceptible to humans but led artificial neural networks to misclassify images, which motivated methods for training artificial neural networks so as to minimize their susceptibility to these adversarial examples (Goodfellow et al., 2014; Madry, 2017). These observations suggest a need for metrics that compare the local geometries of image representations and, in particular, highlight the differences between systems even when global structure seems similar.
@@
We take inspiration from Zhou et al. (2023). For a pair of models and a base image, they synthesize distortions along which the two models’ sensitivities maximally disagree. This bears conceptual similarity to other methods that construct stimuli to optimally distinguish a pair of models (Wang & Simoncelli, 2008; Golan et al., 2020) or cluster neurons by cell type (Burg et al., 2024), and builds on earlier work that examined “eigen-distortions” along which individual models are maximally/minimally sensitive (Berardino et al., 2017). Specifically, Zhou et al. measure the local sensitivity of a model in terms of its Fisher Information matrix (FIM, Fisher, 1925), a classical tool from statistical estimation theory, and choose the pair of “generalized eigen-distortions” that maximize/minimize the ratio of the two models’ sensitivities. Once these image distortions have been computed, they may be added in varying amounts to a base image to determine the level at which they become visible to a human. These measured human sensitivities can then be compared to those of the models, with the goal of identifying which model is better aligned with the local geometry of the human visual system. However, there is no principled method for selecting image distortions for use in comparing more than two models.
@@
We apply our method to a nested set of hand-crafted models of the early visual system to identify distortions that differentiate these models and can potentially be used to evaluate how well these models predict human visual sensitivities. We then apply our method to a set of visual deep neural networks (DNNs) with varying architectures and training procedures. We find distortions that allow for visualization of differences in the sensitivities between layers of the networks and neural network architectures. We further explore differences between standard ImageNet trained networks and their shape-bias enhanced counterparts, and between standard networks and their adversarially-trained counterparts. In all cases, we illustrate how the method generates novel distortions that highlight differences between models.
Portions of this work were initially presented in (Lipshutz et al., 2024).
Figure 1: Comparing the local geometry of image representations. A) Each model maps stimuli to stochastic responses in a model’s representation space—biological neurons are inherently noisy, while a putatively deterministic model can be made stochastic by assuming additive Gaussian response noise. For example, Model A maps the stimulus ${\mathbf{s}}$ (solid black circle $\bullet$) to a conditional density $p({\mathbf{r}}_{A}|{\mathbf{s}})$ on Model A’s representation space (the solid blue circle $\bullet$ and surrounding transparent blue ellipse represent the mean and covariance of ${\mathbf{r}}_{A}$ conditioned on ${\mathbf{s}}$). B) A model’s sensitivity at the base image ${\mathbf{s}}$ to local distortions can be mapped back to the stimulus domain via the model’s positive semidefinite FIM. In the top panel “Eigen-distortions (Berardino et al., 2017)”, the blue ellipse represents the unit level set $\{{\bm{v}}:d_{A}({\bm{v}})=1\}$ of the norm induced by Model A’s FIM ${\bm{I}}_{A}$, which is the set of distortions of the base stimulus ${\mathbf{s}}$ that appear equally distorted according to Model A’s representation. The eigenvectors of the FIM (${\bm{\epsilon}}_{1}$, ${\bm{\epsilon}}_{2}$) can equivalently be interpreted as the distortions that maximize the magnitude of the log ratio of the model’s sensitivities, which is represented as a solid blue circle $\bullet$ on the number line. In the middle panel “Generalized Eigen-distortions (Zhou et al., 2023)”, the blue ellipse is copied from the top panel and the orange ellipse is the level set of Model B’s FIM ${\bm{I}}_{B}$. The generalized eigenvectors of ${\bm{I}}_{A}$ and ${\bm{I}}_{B}$ (${\bm{\epsilon}}_{1}$, ${\bm{\epsilon}}_{2}$) can equivalently be interpreted as the vectors that maximize the difference between the log ratios of the models’ sensitivities. In the bottom panel “Principal distortions (this paper)”, the blue and orange ellipses are as in the above panels and the green ellipse represents the level set of Model C’s FIM ${\bm{I}}_{C}$. We propose a general method for comparing an arbitrary number of models, by selecting two stimulus distortions (${\bm{\epsilon}}_{1}$, ${\bm{\epsilon}}_{2}$) that maximize the variance of the log ratios of model sensitivities.
@@ 2.1 Local information geometry of stochastic image representations
We assume that each image representation has an associated conditional density $p({\mathbf{r}}|{\mathbf{s}})$, where ${\mathbf{s}}$ is a $K$-dimensional vector of image pixels and ${\mathbf{r}}$ is a vector of stochastic responses (e.g., biological neuronal firing rates or deterministic neural network activations with additive response noise). Figure 1A depicts a two-dimensional stimulus space and three models mapping stimuli ${\mathbf{s}}$, ${\mathbf{s}}+{\bm{\epsilon}}_{1}$, and ${\mathbf{s}}+{\bm{\epsilon}}_{2}$ to conditional densities. Note that the dimension of the responses may vary across representations.
@@
where $\nabla_{\mathbf{s}}\log p({\mathbf{r}}|{\mathbf{s}})$ denotes the gradient of $\log p({\mathbf{r}}|{\mathbf{s}})$ with respect to ${\mathbf{s}}$. The FIM is a standard tool in statistical estimation theory that locally approximates the expected log likelihood ratio (or KL divergence) between the conditional distributions $p({\mathbf{r}}|{\mathbf{s}})$ and $p({\mathbf{r}}|{\mathbf{s}}+{\bm{\epsilon}})$, and provides a lower bound on the variance of any unbiased estimator of ${\mathbf{s}}$ (Cramér, 1946; Rao, 1945). The FIM has also been used to link neural representations to perceptual discrimination (Paradiso, 1988; Seung & Sompolinsky, 1993; Brunel & Nadal, 1998; Averbeck & Lee, 2006; Seriès et al., 2009; Ganguli & Simoncelli, 2014; Wei & Stocker, 2016). Given the FIM, we can define the sensitivity of the representation of stimulus ${\mathbf{s}}$ to a distortion ${\bm{\epsilon}}$ as:
@@
As a tractable example, suppose that the conditional response ${\mathbf{r}}$ is Gaussian with stimulus-dependent mean ${\bm{f}}({\mathbf{s}})$ and constant covariance ${\bm{\Sigma}}$. Then the FIM at ${\mathbf{s}}$ is
@@
where ${\bm{J}}_{f}({\mathbf{s}})$ is the Jacobian of ${\bm{f}}(\cdot)$ at ${\mathbf{s}}$. From this expression, we see that the sensitivities of a representation to a distortion ${\bm{\epsilon}}$ depend on how the mean response is changing in the direction ${\bm{\epsilon}}$ relative to the noise covariance.
@@ 2.3 Generalized eigen-distortions for comparing two image representations
${\bm{\epsilon}}_{1}({\mathbf{s}})$ $=\operatorname*{arg\,max}_{\bm{\epsilon}}\frac{{\bm{\epsilon}}^{\top}{\bm{I}}_{A}({\mathbf{s}}){\bm{\epsilon}}}{{\bm{\epsilon}}^{\top}{\bm{I}}_{B}({\mathbf{s}}){\bm{\epsilon}}},$ ${\bm{\epsilon}}_{2}({\mathbf{s}})=\operatorname*{arg\,min}_{\bm{\epsilon}}\frac{{\bm{\epsilon}}^{\top}{\bm{I}}_{A}({\mathbf{s}}){\bm{\epsilon}}}{{\bm{\epsilon}}^{\top}{\bm{I}}_{B}({\mathbf{s}}){\bm{\epsilon}}}.$ (2)
@@ 3 Principal distortions of image representations
We propose a natural extension of generalized eigen-distortions that allows for comparisons among more than two image representations. We show that the generalized eigenvalue problem suggests a metric on image representations, which can be used to optimally choose image distortions that distinguish more than two models.
@@ 3.1 A metric on the local geometry of image representations
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}\left\{\log\frac{{\bm{\epsilon}}^{\top}{\bm{I}}_{A}{\bm{\epsilon}}}{{\bm{\epsilon}}^{\top}{\bm{I}}_{B}{\bm{\epsilon}}}-\log\frac{{\bm{\epsilon}}^{\prime\top}{\bm{I}}_{A}{\bm{\epsilon}}^{\prime}}{{\bm{\epsilon}}^{\prime\top}{\bm{I}}_{B}{\bm{\epsilon}}^{\prime}}\right\},$
where we’ve simplified notation by omitting the dependence of ${\bm{I}}_{*}$ and ${\bm{\epsilon}}_{*}$ on ${\mathbf{s}}$. We can regroup the numerators and denominators of the log quotients by model rather than by distortion, to re-express the optimization problem as follows:
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}({\bm{I}}_{A},{\bm{I}}_{B}),$
@@
$m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}({\bm{I}}_{A},{\bm{I}}_{B}):=\left|\log\frac{d_{A}({\bm{\epsilon}})}{d_{A}({\bm{\epsilon}}^{\prime})}-\log\frac{d_{B}({\bm{\epsilon}})}{d_{B}({\bm{\epsilon}}^{\prime})}\right|$ (3)
and we have used the definition of the sensitivity in Equation 1 and dropped a factor of 2. For any pair of distortions ${\bm{\epsilon}},{\bm{\epsilon}}^{\prime}$, the function $m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}(\cdot,\cdot)$ is a proper metric on positive semi-definite matrices. Specifically, it is non-negative, symmetric, obeys the triangle inequality, and is zero when ${\bm{I}}_{A}={\bm{I}}_{B}$. This metric has several appealing properties:
@@
    $m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}({\bm{I}}_{A},{\bm{I}}_{B})=m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}(c_{A}{\bm{I}}_{A},c_{B}{\bm{I}}_{B}).$
@@
  • Invariance to permutation of the distortions ${\bm{\epsilon}}$ and ${\bm{\epsilon}}^{\prime}$:
    $m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}({\bm{I}}_{A},{\bm{I}}_{B})=m_{{\bm{\epsilon}}^{\prime},{\bm{\epsilon}}}({\bm{I}}_{A},{\bm{I}}_{B}).$
@@
  • The metric compares stochastic representations back in stimulus space via their FIMs. This avoids the problem of aligning the stochastic representations, which is a computationally intensive step that is required when evaluating existing metrics on stochastic representations (Duong et al., 2023b).

### v2

@@ Abstract
Image representations (artificial or biological) are often compared in terms of their global geometric structure; however, representations with similar global structure can have strikingly different local geometries. Here, we propose a framework for comparing a set of image representations in terms of their local geometries. We quantify the local geometry of a representation using the Fisher information matrix, a standard statistical tool for characterizing the sensitivity to local stimulus distortions, and use this as a substrate for a metric on the local geometry in the vicinity of a base image. This metric may then be used to optimally differentiate a set of models, by finding a pair of “principal distortions” that maximize the variance of the models under this metric. As an example, we use this framework to compare a set of simple models of the early visual system, identifying a novel set of image distortions that allow immediate comparison of the models by visual inspection. In a second example, we apply our method to a set of deep neural network models and reveal differences in the local geometry that arise due to architecture and training types. These examples demonstrate how our framework can be used to probe for informative differences in local sensitivities between complex models, and suggest how it could be used to compare model representations with human perception.
@@ 1 Introduction
Biological and artificial neural networks transform sensory stimuli into high-dimensional internal representations that support downstream tasks, and these representations are often described in terms of their neural population geometry (Chung & Abbott, 2021). This idea has led to a multitude of proposed measures of representational similarity (Kriegeskorte et al., 2008; Yamins & DiCarlo, 2016; Kornblith et al., 2019; Williams et al., 2021; Klabunde et al., 2023), which have been used to compare representations within a computational model to the representations within a brain. However, despite differing in architectures and training procedures, many computational models of perceptual or neural responses are equally performant on these representational similarity measures (Schrimpf et al., 2018; Tuckute et al., 2023; Conwell et al., 2024). Are these models functionally interchangeable, or are the datasets and methods that are used to test them simply insufficient to reveal their differences?
Often the similarity between two representations is quantified by measuring alignment of the representations over a set of natural stimuli that are relatively far apart in stimulus space. In this way, these measures capture notions of global geometric similarity between representations. However, systems with similar global structure can have strikingly different local geometries. These local geometries can be investigated by measuring the sensitivity of a system to small image distortions. For example, Szegedy et al. (2013) found image distortions that were imperceptible to humans but led artificial neural networks to misclassify images, which motivated methods for training artificial neural networks so as to minimize their susceptibility to these adversarial examples (Goodfellow et al., 2014; Madry, 2017). These observations suggest a need for metrics that compare the local geometries of image representations and, in particular, highlight the differences between systems even when global structure seems similar. Furthermore, such metrics may also aid in the development of network interpretability and explainability tools.
@@
Figure 1: Comparing the local geometry of image representations. A) Each model is assumed to map stimuli to stochastic responses in a representation space—deterministic models can be investigated by assuming additive Gaussian response noise. For example, Model A maps the stimulus ${\mathbf{s}}$ (solid black circle $\bullet$) to a conditional density $p({\mathbf{r}}_{A}|{\mathbf{s}})$ in Model A’s representation space (solid blue circle $\bullet$ and surrounding transparent blue ellipse). B) Model sensitivity at the base image ${\mathbf{s}}$ to local distortions can be mapped back to the stimulus domain via the model’s positive semidefinite FIM. In the top panel “Eigen-distortions (Berardino et al., 2017)”, the blue ellipse represents the unit level set $\{{\bm{v}}:d_{A}({\bm{v}})=1\}$ of the norm induced by Model A’s FIM ${\bm{I}}_{A}$, which is the set of distortions of the base stimulus ${\mathbf{s}}$ that appear equally distorted according to Model A’s representation. The eigenvectors of the FIM (${\bm{\epsilon}}_{1}$, ${\bm{\epsilon}}_{2}$) can equivalently be interpreted as the distortions that maximize the magnitude of the log ratio of the model’s sensitivities (solid blue circle $\bullet$ on the number line). In the middle panel “Generalized Eigen-distortions (Zhou et al., 2023)”, the blue ellipse is copied from the top panel and the orange ellipse is the level set of Model B’s FIM ${\bm{I}}_{B}$. The generalized eigenvectors of ${\bm{I}}_{A}$ and ${\bm{I}}_{B}$ (${\bm{\epsilon}}_{1}$, ${\bm{\epsilon}}_{2}$) can equivalently be interpreted as the vectors that maximize the difference between the log ratios of the models’ sensitivities. In the bottom panel “Principal distortions (this paper)”, the blue and orange ellipses are as in the above panels and the green ellipse represents the level set of Model C’s FIM ${\bm{I}}_{C}$. Here, we see two stimulus distortions (${\bm{\epsilon}}_{1}$, ${\bm{\epsilon}}_{2}$) that maximize the variance of the log ratios of model sensitivities.
We take inspiration from Zhou et al. (2023). For a pair of models and a base image, they synthesize distortions along which the two models’ sensitivities maximally disagree. This bears conceptual similarity to other methods that construct stimuli to optimally distinguish a pair of models (Wang & Simoncelli, 2008; Golan et al., 2020) or cluster neurons by cell type (Burg et al., 2024), and builds on earlier work that examined “eigen-distortions” along which individual models are maximally/minimally sensitive (Berardino et al., 2017). Specifically, Zhou et al. measure the local sensitivity of a model in terms of its Fisher Information matrix (FIM, Fisher, 1925), a classical tool from statistical estimation theory, and choose the pair of “generalized eigen-distortions” that maximize/minimize the ratio of the two models’ sensitivities. Once these image distortions have been computed, they may be added in varying amounts to a base image to determine the level at which they become visible to a human. These measured human sensitivities can then be compared to those of the models, with the goal of identifying which model is better aligned with the local geometry of the human visual system. The distortions can also be used as model interpretability tools. For instance, visualizations of the distortions may offer insight into how complex models systematically differ, providing practitioners with better understanding of model vulnerabilities. However, there is no principled method for selecting image distortions for comparing more than two models.
@@
We apply our method to a nested set of hand-crafted models of the early visual system to identify distortions that differentiate these models and can potentially be used to evaluate how well these models predict human visual sensitivities. We then apply our method to a set of visual deep neural networks (DNNs) with varying architectures and training procedures. We find distortions that allow for visualization of differences in the sensitivities between layers of the networks and neural network architectures. We further explore differences between standard ImageNet trained networks and their shape-bias enhanced counterparts, and between standard networks and their adversarially-trained counterparts. In all cases, we illustrate how the method generates novel distortions that highlight differences between models. Portions of this work were initially presented in (Lipshutz* et al., 2024).
@@ 2.1 Local information geometry of stochastic image representations
We assume that each image representation has an associated conditional density $p({\mathbf{r}}|{\mathbf{s}})$, where ${\mathbf{s}}$ is a $K$-dimensional vector of image pixels and ${\mathbf{r}}$ is a vector of stochastic responses (e.g., biological neuronal firing rates or deterministic model activations with additive response noise). Fig. 1A depicts a two-dimensional stimulus space and three models mapping stimuli ${\mathbf{s}}$, ${\mathbf{s}}+{\bm{\epsilon}}_{1}$, and ${\mathbf{s}}+{\bm{\epsilon}}_{2}$ to conditional densities. Note that the dimensions of the responses may vary across representations.
@@
where $\nabla_{\mathbf{s}}\log p({\mathbf{r}}|{\mathbf{s}})$ denotes the gradient of $\log p({\mathbf{r}}|{\mathbf{s}})$ with respect to ${\mathbf{s}}$. The FIM is a standard tool in statistical estimation theory that locally approximates the expected log-likelihood ratio (or KL divergence) between the conditional distributions $p({\mathbf{r}}|{\mathbf{s}})$ and $p({\mathbf{r}}|{\mathbf{s}}+{\bm{\epsilon}})$, and provides a lower bound on the variance of any unbiased estimator of ${\mathbf{s}}$ (Cramér, 1946; Rao, 1945). The FIM has also been used to link neural representations to perceptual discrimination (Paradiso, 1988; Seung & Sompolinsky, 1993; Brunel & Nadal, 1998; Averbeck & Lee, 2006; Seriès et al., 2009; Ganguli & Simoncelli, 2014; Wei & Stocker, 2016). Given the FIM, we can define the sensitivity of the representation of stimulus ${\mathbf{s}}$ to a distortion ${\bm{\epsilon}}$ as:
@@
As a tractable example, suppose that the conditional response ${\mathbf{r}}$ is Gaussian with stimulus-dependent mean ${\bm{f}}({\mathbf{s}})$ and constant covariance ${\bm{\Sigma}}$; that is, $p({\mathbf{r}}|{\mathbf{s}})\sim{\mathcal{N}}({\bm{f}}({\mathbf{s}}),{\bm{\Sigma}})$. Then the FIM at ${\mathbf{s}}$ is
@@
where ${\bm{J}}_{f}({\mathbf{s}})$ is the Jacobian of ${\bm{f}}(\cdot)$ at ${\mathbf{s}}$. From this expression, we see that the sensitivities of a representation to a distortion ${\bm{\epsilon}}$ depend on how the mean response is changing in the direction ${\bm{\epsilon}}$ relative to the noise covariance. This example is relevant to our experimental results, where the function ${\bm{f}}({\mathbf{s}})$ denotes the output of a deterministic model evaluated at a stimulus ${\mathbf{s}}$.
@@ 2.3 Generalized eigen-distortions for comparing two image representations
${\bm{\epsilon}}_{1}({\mathbf{s}})$ $=\operatorname*{arg\,max}_{\mathbf{u}}\frac{{\mathbf{u}}^{\top}{\bm{I}}_{A}({\mathbf{s}}){\mathbf{u}}}{{\mathbf{u}}^{\top}{\bm{I}}_{B}({\mathbf{s}}){\mathbf{u}}},$ ${\bm{\epsilon}}_{2}({\mathbf{s}})=\operatorname*{arg\,min}_{\mathbf{v}}\frac{{\mathbf{v}}^{\top}{\bm{I}}_{A}({\mathbf{s}}){\mathbf{v}}}{{\mathbf{v}}^{\top}{\bm{I}}_{B}({\mathbf{s}}){\mathbf{v}}}.$ (2)
@@ 3 Principal distortions of image representations
We propose a natural extension of generalized eigen-distortions that allows for comparisons among more than two image representations. We show that the generalized eigenvalue problem suggests a metric on the local geometry of image representations, which can be used to optimally choose image distortions that distinguish more than two models at a base stimulus.
@@ 3.1 A metric on the local geometry of image representations
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\mathbf{u}},{\mathbf{v}}}\left|\log\frac{d_{A}({\mathbf{u}})}{d_{B}({\mathbf{u}})}-\log\frac{d_{A}({\mathbf{v}})}{d_{B}({\mathbf{v}})}\right|.$
where we’ve used the definition of the sensitivity in equation 1 and simplified notation by omitting the dependence of sensitivities and distortions on the stimulus, ${\mathbf{s}}$. We can regroup the numerators and denominators of the log quotients by model rather than by distortion, to re-express the optimization problem as follows:
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\mathbf{u}},{\mathbf{v}}}m_{{\mathbf{u}},{\mathbf{v}}}({\bm{I}}_{A},{\bm{I}}_{B}),$
@@
$m_{{\mathbf{u}},{\mathbf{v}}}({\bm{I}}_{A},{\bm{I}}_{B}):=\left|\log\frac{d_{A}({\mathbf{u}})}{d_{A}({\mathbf{v}})}-\log\frac{d_{B}({\mathbf{u}})}{d_{B}({\mathbf{v}})}\right|.$ (3)
For any pair of distortions ${\mathbf{u}},{\mathbf{v}}$, the function $m_{{\mathbf{u}},{\mathbf{v}}}(\cdot,\cdot)$ defines a (pseudo-)metric. [^1] Technically, $m_{{\mathbf{u}},{\mathbf{v}}}(\cdot,\cdot)$ is a pseudometric because $m_{{\mathbf{u}},{\mathbf{v}}}({\bm{I}}_{A},{\bm{I}}_{B})=0$ does not imply that ${\bm{I}}_{A}={\bm{I}}_{B}$. Specifically, it is non-negative, symmetric, obeys the triangle inequality, and is zero when ${\bm{I}}_{A}={\bm{I}}_{B}$. This metric has several appealing properties:
@@
    $m_{{\mathbf{u}},{\mathbf{v}}}({\bm{I}}_{A},{\bm{I}}_{B})=m_{{\mathbf{u}},{\mathbf{v}}}(c_{A}{\bm{I}}_{A},c_{B}{\bm{I}}_{B}).$
@@
  • Invariance to permutation of the distortions ${\mathbf{u}}$ and ${\mathbf{v}}$:
    $m_{{\mathbf{u}},{\mathbf{v}}}({\bm{I}}_{A},{\bm{I}}_{B})=m_{{\mathbf{u}},{\mathbf{v}}}({\bm{I}}_{A},{\bm{I}}_{B}).$
@@
  • The metric compares stochastic representations back in stimulus space via their FIMs, which avoids having to align the stochastic representations (Duong et al., 2023b).